# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ VIII-A3 Byzantine Attacks Defense Mechanisms
We organize the works that belong to this class into two subclasses, namely architectural style, and aggregation rules, as provided below:
  • Architectural Style: such mechanisms exploit the architectural learning style to detect and eliminate malicious updates. Gu et al. [252] proposed an unsupervised Conditional Variational Autoencoder (CVAE)-based training anomaly detection system, called Fedcvae, which aims to accurately discover and eradicate malicious model updates to negate their harmful consequences. The architecture of Fedcvae consists of an encoder-decoder design. Using the encoder, the original variables are mapped into low-dimensional embeddings, while the decoder rebuilds the initial variables from them. The performance evaluation under four different datasets, namely Vehicle, Synthetic, MNIST, and FEMINIST (under both IID and non-IID data distribution approaches) shows that the proposed system can withstand Byzantine attacks and targeted model poisoning attacks with up to 30% of the participating clients being malicious.
  • Aggregation Rules: The main purpose of these mechanisms is to ensure safe learning, whether all clients are honest or there is a subset of malicious clients. Blanchard et al. [243] investigated Byzantine failure robustness for distributed Stochastic Gradient Descent (SGD)-based frameworks. The research seeks to address the problem of Byzantine failure resilience of an SGD system with no limitation on the dimension or parameter space size. The proposed solution entails the formulation of a resilience property of the aggregation rule that grasps the core demands for ensuring convergence regardless of having a subset of Byzantine parties. This technique chooses and eliminates the most distant vectors with respect to the calculated average from the aggregated gradient updates. The distance is computed using the Euclidean spacing among the gradient vectors. The approach works well against attacks involving up to 33% of adversely affected parties.
@@ VIII-A4 Sybil Attacks Defense Mechanisms
  • Verifiable Random Functions: in short, a verifiable random function (VRF) is given two inputs (a secret key and a seed) and returns two associated values (hash and proof). The proof allows everyone who is carrying the public key of the given peer to confirm that the given hash was indeed produced by a peer who is in possession of the private key. Shayan et al. [254] present Biscotti, a secure blockchain-based framework for fully decentralized multi-party privacy-preserving ML. The proposed system implements a robust hash protocol built upon the most recent block hash and VRFs for subgroup selection among peers who are in the process of carrying out individual steps (noise addition, updates validations, and secure aggregations). The proposed system is based on the Proof of Federation (PoF) protocol, a proposed blockchain consensus protocol that employs FL defenses and then renders them enforceable across decentralized P2P environments. As the selection of subgroups is performed by the VRF, the stake of the peer is treated as the individual’s reputation acquired by contributing favorably to the shared model. This guarantees that an opponent is unable to expand its leverage in the system by establishing too many peers and still not enhancing the model, which is useful for mitigating the effect of Sybils. The experimental results with 26 committee sizes provide immunity against an opponent who controls 30% of the system’s stakes.
  • Contribution Similarity: Fung et al. [147] present a defensive discriminant technique to distinguish Sybils from legitimate users based on the diversity of their relative gradient updates. The proposed learning parameters make use of a per-client adaptive learning rate that relies on the contribution similarity among the clients. The identification of malicious users is based on a common goal, and therefore their model updates are susceptible to low variance, which is done by searching for similarities regarding indicative characteristics. Experimental evaluations on four datasets, namely MNIST, VGGFace2, KDDCup, and Amazon, with 5 attack scenarios indicated that the proposed system, in conjunction with other modules (such as Multi-Krum), can successfully mitigate a range of different attack types and even when Sybils submerge legitimate users.
@@ VIII-A5 Defense Mechanisms Against Combined Attacks
We organize the works that belong to this class into two subclasses, including Moving Target Defense and Asynchronous Convergence.
  • Moving Target Defense: the purpose of MTD is to actively shift the surface available for attacks. Qiu et al. [180] propose MT-MTD, an MTD-based defense strategy against Trojan attacks on DNN networks. The framework takes an attack-defense game approach, specifically, a signaling game. MT-MTD has four main steps. The first is a dimensional division of the training set by the defender and then passed to the attacker for training. The second is a random selection from the resulting collection of dimensional combinations. The third step is the weight adjustment. The last is the consensus process.
  • Asynchronous Convergence: used in situations when the aggregation can be done asynchronously. Liu et al. [253] propose a security architecture dedicated to a secure FL-based learning workflow. The architecture is composed of two parts. The first one, named FedBlock, is proposed to introduce decentralization in the learning process through the use of the blockchain, while the second one, named FedAC, permits the FL to carry out a global aggregation in an asynchronous manner while considering a staleness coefficient. The system is dedicated to securing the learning process against a multitude of threats, including Single Point Of Failure (SPOF), unconventional learning failures, and dedicated attacks. Experimental evaluation on a physical deployment, namely a Raspberry Pi (4b) under the MNIST dataset, shows that the system can reach accuracy scores as high as 98.96% in horizontal FL and 95.84% in vertical FL. The reported results of experimental evaluations on the CIFAR-10 and Fashion-MNIST datasets show that the proposed approach can be effective in identifying and differentiating honest updates from malicious ones.
@@ VIII-B1 Trained Model Integrity Verification
  • Cryptographically-protected Provenance: these defensive mechanisms protect the model trained using encryption. Unal et al. [263] propose a security scheme based on Blockchain and computing fuzzy hash values for protecting FL algorithms operating in IoT systems. Notably, the proposed scheme stores the model parameters in the blockchain, although this is true to some extent, the model parameters on the blockchain are not stored unchanged. In other words, the authors applied a one-way hash function upon these model parameters prior to their storage on the blockchain. Thus, the introduced scheme provides an efficient approach that does not affect FL’s privacy concerns. In another approach under the same context, Stokes et al. [249] propose a cryptography-based authentication, which provides a cryptography-based provenance module to protect trained models against alteration of their settings or underlying structure. The proposed authentication can protect against model poisoning attacks.
  • Optimization-based: in a work by Kuttichira et al. [264], the authors propose a Bayesian Compromise Detection (BCD) algorithm that aims to address the potential security risk associated with malignant modification of stored models in the cloud. The main solution consists of solving an optimization problem that seeks to maximize the discrepancies in prediction outcomes by comparing the genuine model with the damaged model. Although this task seems easy at first glance, the difficulty of the problem lies in three different points according to the authors. First, its non-convex nature, second, the large space dimensionality when searching through the training input distribution for the sensitive sample, and third, a black-box view of the compromised model is all that cloud clients can have. The proposed solution consists of two parts. First, a Variational Autoencoder (VAE) is used to associate high-dimensional data with a low-dimensional nonlinear space, and second, Bayesian optimization (BO) is applied to determine the generally ideal sensitive sample, which can identify the model corruption with little overhead. Experiments on MNIST, Olivetti, and CIFAR-100 datasets illustrate the capability of the proposed approach, with results of up to 100% detection rate.
@@ VIII-C1 Privacy Leakage Defense Mechanisms
  • Gradients Shielding: the main objective of these mechanisms is to protect the gradient of the model in a cryptographic way. Hao et al. [265] implemented a privacy-preserving IDS for FL-based industrial environments, named PEFL. The proposed system secures exchanged aggregation with homomorphic encryption and differential privacy. In essence, the homomorphically encrypted data of the private gradients is incorporated into the A-LWE (augmented learning with error). Benchmarking on the MNIST dataset exhibits good performance, along with computational and communication cost efficiency.
  • Self-Distillation: in a work by Tang et al. [262], the authors propose SELENA, a privacy-preserving model learning framework that generates equivalent member/non-member feedback to alleviate black-box membership inference threats. The proposed system consists of two major components. The first is the Split Adaptive Inference Ensemble (Split-AI), which allows the model to behave similarly to members and non-members samples. This is done by training sub-models through random subsets within the training set. The second is a self-distillation mechanism that transfers knowledge about the model created by Split-AI to deliver a protected end model. To do this, Split-AI is first interrogated with the exact training data to retrieve the related prediction sequences. Then, with these predictions as soft labels, the protected end model is trained. Evaluations on three different datasets (Purchase100, Texas100, and CIFAR100), with ResNet-18 and a fully connected four-layer NN, and different MIA attacks, including direct single-query attacks, label-only attacks, and adaptive attacks, showed that the proposed system achieved a good tradeoff between practicality and privacy.
  • Overfitting Control: when given a data point on which models have been trained, they return a high aftereffect value on a class relative to the others, reflecting the underlying overfitting nature of ML models, which is taken to be one of the reasons for the effectiveness of MIAs. Hence, according to Salem et al. [266] controlling overfitting is one way to mitigate such attacks. The authors propose the use of two approaches: one classical and the other relatively new, namely dropout and model stacking. The first one is a regularization technique that prevents complex co-matching on training data and is specifically used for DL models. The second one is proposed to be used for other ML classifiers. The main concept is to hierarchically organize several ML models in order to avoid overfitting.
@@ IX Lessons learned, Open issues and Challenges
Despite the significant efforts made by the scientific community to strengthen security in the cyber world in general, and for future networks such as 6G, in collaboration with emerging technologies such as IoT and AI in particular, there is however a long way to be traveled, if we are to achieve a fully secure cyber environment. In this section, we present the lessons learned, open issues, and challenges.
@@ IX-A Lessons learned
Through comprehensive reviews and in-depth analysis, we were able to classify the datasets used by the scientific community for experimenting and evaluating ML techniques on cyber attacks into seven main categories, including, IoT-based multi-purpose security, attack classification, image classification, time series classification, human activity recognition, sentiment classification, location awareness, and text classification. From the attacks against machine learning systems, we found twenty-one attacks. According to the actual context of the attack in edge learning, we were able to classify them into eight categories, including, backdoor attacks, adversarial examples, combined attacks, poisoning attacks, Sybil attacks, byzantine attacks, inference attacks, and drop attacks. Based on the deployment strategy of each security defense, we were able to classify the defense methods against edge learning vulnerabilities into three categories, including, training phase defense methods, post-training phase defense methods, and inference phase defense methods.
@@
  • Definition of the infrastructure of the 6G-enabled IoT network as well as the emerging technologies adopted for each layer (e.g., Edge layer, Fog Layer, SDN Layer, Blockchain Layer, Digital twins layer…etc).
  • Definition of the edge learning model (i.e., centralized learning, federated learning, distributed learning).
@@
  • Experimenting and evaluating the defense method in terms of accuracy, precision, recall, and attack rate.
  • Study the performances of the system with the application of the proposed defense method in terms of scalability and interoperability.
The use of network slicing, multi-access edge computing, cloud computing, virtualization, terahertz and sub-terahertz communications, and artificial intelligence in the 5G/6G IoT testbeds demonstrates the complexity and diversity of the technologies being developed. These testbeds offer an opportunity to explore new possibilities in smart transportation, energy management, healthcare, and industrial automation.
@@
TABLE XVII: Open Issues and Challenges in 6G-IoT edge learning vulnerabilities.
@@ IX-B Open issues and Challenges
Table XVII provides an overview of the open issues and challenges in 6G-IoT machine learning vulnerabilities. The challenges covered include reliable and trustworthy learning for 6G-IoT intelligence, security solutions adaptability, ethics by design, datasets formation/availability, learning complexity, high-quality data availability, ML vulnerabilities elimination, and defense strategies implementations.
@@ IX-B1 Reliable and Trustworthy Learning for 6G-IoT Intelligence
AI capabilities have proven to be a key component of future technologies, however, there is more involved in ensuring security. Here we present some of the challenges to be addressed when creating reliable and trustworthy learning environments.
@@
  • Natively Secured AI: Adversary-aware AI, or natively secured AI, is an upcoming class for AI-based systems that focuses on building security into the data preparation, learning, storage, and inference stages. This paradigm is necessary because cyberattacks can compromise the integrity, confidentiality, and availability of AI models and their associated data, leading to adverse consequences. Similar to how cyber attacks forced developers to write secure code and libraries for programming languages, the increasing adoption of AI-based systems will necessitate the adoption of secure AI practices. This will involve not only the development of secure AI algorithms but also the design and implementation of secure hardware and software architectures to support these algorithms. The challenges associated with implementing natively secured AI are considerable, given the complexity and uncertainty associated with AI techniques. Additionally, the limitations of IoT devices, such as their limited processing and storage resources, may pose a challenge to running complex AI models. To address this challenge, lightweight and efficient ML algorithms tailored for these devices must be developed, similar to the situation encountered in the case of lightweight IoT-based authentication protocols. In summary, natively secured AI is an upcoming class for AI-based systems that focuses on building security into every aspect of AI, from data preparation to inference. While there are significant challenges associated with implementing this paradigm, it is necessary to ensure the security of AI-based systems in the face of increasingly sophisticated cyber attacks.
@@
  • Trustworthy AI: Trustworthy AI is a key component of natively secure AI, assuring that AI-based systems operate reliably, transparently, and ethically. Specifically, Trustworthy AI refers to AI that is designed to be transparent, explainable, and accountable. Trustworthy AI systems focus on minimizing the risks associated with AI algorithms and their decision-making processes by ensuring that they operate within specific ethical constraints. To achieve trustworthy AI, secure AI algorithms must be developed that incorporate ethical and legal standards, including human rights, data protection, and privacy. This is essential to ensure that AI-based systems are transparent and explainable and that users can understand how the system works, how it makes decisions, and how it uses data.
@@
  • Security Solutions Adaptability: The dynamic nature of settings shifts across wireless networks necessitates that the proposed security solutions must be highly adaptive. Due to the fact that many operators collaborate with each other, this diversity can pose a significant heterogeneity problem if not handled effectively. Solutions to this problem have been briefly discussed, including pre-trained models, unified datasets, and agreed-upon models between Next Generation NodeB (gNB) and user equipment (UEs) [267]. However, real-world implementations and further tests are different stories with varying perspectives, posing a future challenge for these networks. Hence, the adaptability of security solutions is crucial for the security of wireless networks, and ongoing research and testing are necessary to address the challenges posed by the dynamic and heterogeneous nature of these networks.
@@
  • Ethics by Design: While machine intelligence is a promising addition to future networks, making them fully automated, they do not accommodate ethical guidelines in the same way as humans [5]. The "Ethics by Design" approach addresses the ethical dimensions of AI systems at the earliest stage of their design [268]. And as much as it is not an easy subject for a machine to learn, the concept itself is difficult to define, as even among humans and nations, the concept sometimes has wide variations in how it is described. By adopting the Ethics by Design approach, AI designers and developers can help ensure that AI systems are developed in a responsible and ethical manner.

### v2

@@ VIII-A3 Byzantine Attacks Defense Mechanisms
We organize the works that belong to this class into two subclasses, namely architectural style and aggregation rules, as provided below:
  • Architectural Style: Such mechanisms exploit the architectural learning style to detect and eliminate malicious updates. Gu et al. [279] proposed an unsupervised Conditional Variational Autoencoder (CVAE)-based training anomaly detection system, called Fedcvae, which aims to accurately discover and eradicate malicious model updates to negate their harmful consequences. The architecture of Fedcvae consists of an encoder-decoder design. Using the encoder, the original variables are mapped into low-dimensional embeddings while the decoder rebuilds their initial variables. The performance evaluation under four different datasets, namely Vehicle, Synthetic, MNIST, and FEMINIST (under both IID and non-IID data distribution approaches) shows that the proposed system can withstand Byzantine attacks and targeted model poisoning attacks with up to 30% of the participating clients being malicious.
  • Aggregation Rules: The main purpose of these mechanisms is to ensure safe learning, whether all clients are honest or there is a subset of malicious clients. Blanchard et al. [270] investigated Byzantine failure robustness for distributed Stochastic Gradient Descent (SGD)-based frameworks. The research seeks to address the problem of Byzantine failure resilience of an SGD system with no limitation on the dimension or parameter space size. The proposed solution entails formulating a resilience property of the aggregation rule that grasps the core demands for ensuring convergence regardless of having a subset of Byzantine parties. This technique chooses and eliminates the most distant vectors concerning the calculated average from the aggregated gradient updates. The distance is computed using the Euclidean spacing among the gradient vectors. The approach works well against attacks involving up to 33% of adversely affected parties.
@@ VIII-A4 Sybil Attacks Defense Mechanisms
  • Verifiable Random Functions: In short, a verifiable random function (VRF) is given two inputs (a secret key and a seed) and returns two associated values (hash and proof). The proof allows everyone who is carrying the public key of the given peer to confirm that the given hash was indeed produced by a peer who has the private key. Shayan et al. [281] present Biscotti, a secure blockchain-based framework for fully decentralized multi-party privacy-preserving ML. The proposed system implements a robust protocol built upon the most recent block hash and VRFs for subgroup selection among peers carrying out individual steps (noise addition, updates validations, and secure aggregations). The proposed system is based on the Proof of Federation (PoF) protocol, a blockchain consensus protocol that employs FL defenses and renders them enforceable across decentralized P2P environments. As the selection of subgroups is performed by the VRF, the stake of the peer is treated as the individual’s reputation acquired by contributing favorably to the shared model. This guarantees that an opponent cannot expand its leverage in the system by establishing too many peers and still not enhancing the model, which helps mitigate the effect of Sybils. The experimental results with 26 committee sizes provide immunity against an opponent who controls 30% of the system’s stakes.
  • Contribution Similarity: Fung et al. [168] present a defensive discriminant technique to distinguish Sybils from legitimate users based on the diversity of their relative gradient updates. The proposed learning parameters use a per-client adaptive learning rate that relies on the contribution similarity among the clients. The identification of malicious users is based on a common goal. Therefore their model updates are susceptible to low variance, which is done by searching for similarities regarding indicative characteristics. Experimental evaluations on four datasets, namely MNIST, VGGFace2, KDDCup, and Amazon, with 5 attack scenarios indicated that the proposed system, in conjunction with other modules (such as Multi-Krum), can successfully mitigate a range of different attack types and even when Sybils submerge legitimate users.
@@ VIII-A5 Defense Mechanisms Against Combined Attacks
We organize the works in this class into two subclasses, including Moving Target Defense and Asynchronous Convergence.
  • Moving Target Defense: MTD aims to shift the surface available for attacks actively. Qiu et al. [200] propose MT-MTD, an MTD-based defense strategy against Trojan attacks on DNN networks. The framework takes an attack-defense game approach, specifically, a signaling game. MT-MTD has four main steps. The first is a dimensional division of the training set by the defender and then passed to the attacker for training. The second is a random selection from the resulting collection of dimensional combinations. The third step is the weight adjustment. The last is the consensus process.
  • Asynchronous Convergence: used when aggregation can be done asynchronously. Liu et al. [280] propose a security architecture dedicated to a secure FL-based learning workflow. The architecture is composed of two parts. The first one, named FedBlock, is proposed to introduce decentralization in the learning process through the use of the blockchain, while the second one, named FedAC, permits the FL to carry out a global aggregation in an asynchronous manner while considering a staleness coefficient. The system is dedicated to securing the learning process against many threats, including Single Point Of Failure (SPOF), unconventional learning failures, and dedicated attacks. Experimental evaluation on a physical deployment, namely a Raspberry Pi (4b) under the MNIST dataset, shows that the system can reach accuracy scores as high as 98.96% in horizontal FL and 95.84% in vertical FL. The reported results of experimental evaluations on the CIFAR-10 and Fashion-MNIST datasets show that the proposed approach can be effective in identifying and differentiating honest updates from malicious ones.
@@ VIII-B1 Trained Model Integrity Verification
  • Cryptographically-protected Provenance: these defensive mechanisms protect the model trained using encryption. Unal et al. [290] propose a security scheme based on Blockchain and computing fuzzy hash values for protecting FL algorithms operating in IoT systems. Notably, the proposed scheme stores the model parameters in the blockchain, although this is true to some extent, the model parameters on the blockchain are not stored unchanged. In other words, the authors applied a one-way hash function upon these model parameters before their storage on the blockchain. Thus, the introduced scheme provides an efficient approach that does not affect FL’s privacy concerns. In another approach under the same context, Stokes et al. [276] propose a cryptography-based authentication, which provides a cryptography-based provenance module to protect trained models against alteration of their settings or underlying structure. The proposed authentication can protect against model poisoning attacks.
  • Optimization-based: in a work by Kuttichira et al. [291], the authors propose a Bayesian Compromise Detection (BCD) algorithm that aims to address the potential security risk associated with malignant modification of stored models in the cloud. The main solution consists of solving an optimization problem that seeks to maximize the discrepancies in prediction outcomes by comparing the genuine and damaged models. Although this task seems easy at first glance, the problem’s difficulty lies in three points, according to the authors. First, its non-convex nature; second, the large space dimensionality when searching through the training input distribution for the sensitive sample; and third, a blue-box view of the compromised model is all that cloud clients can have. The proposed solution consists of two parts. First, a Variational Autoencoder (VAE) is used to associate high-dimensional data with a low-dimensional nonlinear space, and second, Bayesian optimization (BO) is applied to determine the generally ideal sensitive sample, which can identify the model corruption with little overhead. Experiments on MNIST, Olivetti, and CIFAR-100 datasets illustrate the capability of the proposed approach, with results of up to 100% detection rate.
@@ VIII-C1 Privacy Leakage Defense Mechanisms
  • Gradients Shielding: The main objective of these mechanisms is to protect the model’s gradient cryptographically. Hao et al. [292] implemented a privacy-preserving IDS for FL-based industrial environments named PEFL. The proposed system secures exchanged aggregation with homomorphic encryption and differential privacy. In essence, the homomorphically encrypted data of the private gradients is incorporated into the A-LWE (augmented learning with error). Benchmarking on the MNIST dataset exhibits good performance, along with computational and communication cost efficiency.
  • Self-Distillation: in a work by Tang et al. [289], the authors propose SELENA. This privacy-preserving model learning framework generates equivalent member/non-member feedback to alleviate blue-box membership inference threats. The proposed system consists of two major components. The first is the Split Adaptive Inference Ensemble (Split-AI), which allows the model to behave similarly to members and non-members samples. This is done by training sub-models through random subsets within the training set. The second is a self-distillation mechanism that transfers knowledge about the model created by Split-AI to deliver a protected end model. To do this, Split-AI is first interrogated with the exact training data to retrieve the related prediction sequences. Then, the protected end model is trained with these predictions as soft labels. Evaluations on three different datasets (Purchase100, Texas100, and CIFAR100), with ResNet-18 and a fully connected four-layer NN, and different MIA attacks, including direct single-query attacks, label-only attacks, and adaptive attacks, showed that the proposed system achieved a good tradeoff between practicality and privacy.
  • Overfitting Control: when given a data point on which models have been trained, they return a high aftereffect value on a class relative to the others, reflecting the underlying overfitting nature of ML models, which is taken to be one of the reasons for the effectiveness of MIAs. Hence, according to Salem et al. [293] controlling overfitting is one way to mitigate such attacks. The authors propose using two approaches: classical and relatively new, namely dropout and model stacking. The first is a regularization technique that prevents complex co-matching on training data and is used explicitly for DL models. The second one is proposed to be used for other ML classifiers. The main concept is to organize several ML models to avoid overfitting hierarchically.
@@ IX Lessons learned, Open issues and Challenges
Despite the significant efforts made by the scientific community to strengthen security in the cyber world in general, and for future networks such as 6G, in collaboration with emerging technologies such as IoT and AI in particular, there is however a long way to be traveled, if we are to achieve a fully secure cyber environment. This section presents the lessons learned, open issues, and challenges.
@@ IX-A Lessons learned
Through comprehensive reviews and in-depth analysis, we were able to classify the datasets used by the scientific community for experimenting and evaluating ML techniques on cyber attacks into nine main categories, including, IoT-based multi-purpose security, attack classification, image classification, time series classification, human activity recognition, sentiment classification, location awareness, and text classification. From the attacks against machine learning systems, we found twenty-one attacks. According to the actual context of the attack in Edge Learning, we were able to classify them into eight categories, including backdoor attacks, adversarial examples, combined attacks, poisoning attacks, Sybil attacks, byzantine attacks, inference attacks, and drop attacks. Based on the deployment strategy of each security defense, we classified the defense methods against Edge Learning vulnerabilities into three categories, including training phase defense methods, post-training phase defense methods, and inference phase defense methods.
@@
  • Definition of the infrastructure of the 6G-enabled IoT network and the emerging technologies adopted for each layer (e.g., Edge layer, Fog Layer, SDN Layer, Blockchain Layer, Digital twins layer…etc.).
  • Definition of the edge learning model (i.e., centralized, federated, distributed learning).
@@
  • Experimenting and evaluating the defense method regarding accuracy, precision, recall, and attack rate.
  • Study the system’s performances with the application of the proposed defense method regarding scalability and interoperability.
The use of network slicing, multi-access edge computing, cloud computing, virtualization, terahertz, and sub-terahertz communications, and artificial intelligence in the 5G/6G IoT testbeds demonstrates the complexity and diversity of the technologies being developed. These testbeds offer an opportunity to explore new possibilities in smart transportation, energy management, healthcare, and industrial automation.
@@
TABLE XIX: Open Issues and Challenges in 6G-IoT Edge Learning vulnerabilities.
@@ IX-B Open issues and Challenges
Table XIX overviews the open issues and challenges in 6G-IoT machine learning vulnerabilities. The challenges covered include reliable and trustworthy learning for 6G-IoT intelligence, security solutions adaptability, ethics by design, datasets formation/availability, learning complexity, high-quality data availability, ML vulnerabilities elimination, and defense strategies implementations.
@@ IX-B1 Reliable and Trustworthy Learning for 6G-IoT Intelligence
AI capabilities have proven to be a key component of future technologies. However, there is more involved in ensuring security. Here we present some challenges when creating reliable and trustworthy learning environments.
@@
  • Natively Secured AI: Adversary-aware AI, or natively secured AI, is an upcoming class for AI-based systems that focuses on building security into the data preparation, learning, storage, and inference stages. This paradigm is necessary because cyberattacks can compromise the integrity, confidentiality, and availability of AI models and their associated data, leading to adverse consequences. Similar to how cyber attacks forced developers to write secure code and libraries for programming languages, the increasing adoption of AI-based systems will necessitate the adoption of secure AI practices. This will involve not only the development of secure AI algorithms but also the design and implementation of secure hardware and software architectures to support these algorithms. Given the complexity and uncertainty associated with AI techniques, the challenges associated with implementing natively secured AI are considerable. Additionally, the limitations of IoT devices, such as their limited processing and storage resources, may pose a challenge to running complex AI models. To address this challenge, lightweight and efficient ML algorithms tailored for these devices must be developed, similar to the situation encountered in the case of lightweight IoT-based authentication protocols. In summary, natively secured AI is an upcoming class for AI-based systems that focuses on building security into every aspect of AI, from data preparation to inference. While significant challenges are associated with implementing this paradigm, ensuring the security of AI-based systems in the face of increasingly sophisticated cyber attacks is necessary.
@@
  • Trustworthy AI: Trustworthy AI is a key component of natively secure AI, assuring that AI-based systems operate reliably, transparently, and ethically. Specifically, Trustworthy AI refers to AI designed to be transparent, explainable, and accountable. Trustworthy AI systems focus on minimizing the risks associated with AI algorithms and their decision-making processes by ensuring that they operate within specific ethical constraints. To achieve trustworthy AI, secure AI algorithms must be developed that incorporate ethical and legal standards, including human rights, data protection, and privacy. This is essential to ensure that AI-based systems are transparent and explainable and that users can understand how the system works, makes decisions and uses data.
@@
  • Security Solutions Adaptability: The dynamic nature of settings shifts across wireless networks requires the proposed security solutions to be highly adaptive. Because many operators collaborate, this diversity can pose a significant heterogeneity problem if not handled effectively. Solutions to this problem have been briefly discussed, including pre-trained models, unified datasets, and agreed-upon models between Next Generation NodeB (gNB) and user equipment (UEs) [294]. However, real-world implementations and further tests are different stories with varying perspectives, posing a future challenge for these networks. Hence, the adaptability of security solutions is crucial for the security of wireless networks, and ongoing research and testing are necessary to address the challenges posed by these networks’ dynamic and heterogeneous nature.
@@
  • Ethics by Design: While machine intelligence is a promising addition to future networks, making them fully automated, they do not accommodate ethical guidelines in the same way as humans [8]. The "Ethics by Design" approach addresses the ethical dimensions of AI systems at the earliest stage of their design [295]. And as much as it is not an easy subject for a machine to learn, the concept itself is difficult to define, as even among humans and nations, the concept sometimes has wide variations in how it is described. By adopting the Ethics by Design approach, AI designers and developers can help ensure that AI systems are developed responsibly and ethically.